OpenStudy (anonymous):

Given an object in a straight line with a(t)=√t ft/sec^2 and v(0)=-18 for t=0 to t=16, find v(t) and the displacement and distance of the object.

OpenStudy (phi):

this sounds like an integration problem

OpenStudy (anonymous):

It is

OpenStudy (anonymous):

What's the thought process of the problem?

OpenStudy (phi):

velocity is \( \frac{dx}{dt} \) change in distance over change in time acceleration is \( \frac{dv}{dt} = \frac{d^2x}{dt^2} \) the second derivative of x with respect to time \[v= \int\limits_{0}^{16} \frac{dv}{dt} dt= \int\limits_{0}^{16} t^{\frac{1}{2}} dt\]